# Iterative Distillation for Reward-Guided Fine-Tuning of Diffusion Models in Biomolecular Design

### v1

@@ Abstract
We address the problem of fine-tuning diffusion models for reward-guided generation in biomolecular design. While diffusion models have proven highly effective in modeling complex, high-dimensional data distributions, real-world applications often demand more than high-fidelity generation, requiring optimization with respect to potentially non-differentiable reward functions such as physics-based simulation or rewards based on scientific knowledge. Although RL methods have been explored to fine-tune diffusion models for such objectives, they often suffer from instability, low sample efficiency, and mode collapse due to their on-policy nature. In this work, we propose an iterative distillation-based fine-tuning framework that enables diffusion models to optimize for arbitrary reward functions. Our method casts the problem as policy distillation: it collects off-policy data during the roll-in phase, simulates reward-based soft-optimal policies during roll-out, and updates the model by minimizing the KL divergence between the simulated soft-optimal policy and the current model policy. Our off-policy formulation, combined with KL divergence minimization, enhances training stability and sample efficiency compared to existing RL-based methods. Empirical results demonstrate the effectiveness and superior reward optimization of our approach across diverse tasks in protein, small molecule, and regulatory DNA design.
@@ 1 Introduction
Our contributions are summarized as follows. We propose a novel algorithm, VIDD, for fine-tuning diffusion models through iterative distillation of target policies composed of both value functions and the current policy. Unlike direct reward backpropagation methods, which require differentiable reward signals, our approach can handle non-differentiable rewards commonly encountered in scientific domains. Compared to PPO-based methods, our algorithm achieves more stable optimization. Algorithmically, our key innovation lies in effectively incorporating value functions specifically tailored to the diffusion models. Empirically, we validate the effectiveness of our method across a range of scientific tasks, particularly in molecular design.
@@ 1.1 Related Works
An alternative line of work focuses on training-free techniques that aim to improve rewards solely at inference time. The most straightforward example is Best-of-N sampling, which selects the highest-reward output from N generated candidates. More sophisticated methods go further by evaluating and selecting promising intermediate generation states during the sampling process (Wu et al., 2023; Li et al., 2024; Kim et al., 2025; Singhal et al., 2025; Ma et al., 2025). While these methods avoid the need for fine-tuning, they often incur significantly higher inference-time costs due to repeated sampling. Moreover, they typically struggle to surpass the reward performance of pre-trained models by a large margin. In this respect, training-free methods are orthogonal and complementary to fine-tuning-based approaches.
@@ 3.2 Iterative Distillation
$\mathop{\mathrm{argmin}}_{\theta}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}[\mathrm{KL}(p^{\star}_{t-1}(\cdot|x_{t})\|p^{\theta}_{t-1}(\cdot|x_{t}))]$ (4)
@@
$\mathop{\mathrm{argmax}}_{\theta}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\frac{1}{\exp(v_{t}(x_{t})/\alpha)}\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t}))]\right]$ (5)
@@
In practice, we adopt an iterative distillation procedure. This is motivated by the fact that, due to the reliance on empirical samples and approximate value functions, it is generally infeasible to accurately distill the target policy $p^{\star}_{t-1}$ into the student policy $p_{t-1}^{\theta}$ in a single step. Rather than fixing $p^{\mathrm{pre}}_{t-1}$ throughout training, we periodically update it by the latest student policy but in a lazy manner, allowing the target to gradually improve over time. This dynamic yet infrequent (lazy) target update enables progressive refinement, allowing both the teacher and student policies to evolve iteratively toward better alignment. This strategy is analogous to standard RL practices grounded in the Policy Improvement Theorem.
@@
$\theta_{s+1}\leftarrow\theta_{s}+\nabla\sum_{t}\mathbb{E}_{x_{t}\sim u_{t},p^{\mathrm{out}}_{t-1}(x_{t-1}|x_{t})}[\exp(v^{\mathrm{out}}_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t})]$ (6)
where $\{p^{\mathrm{out}}_{t}\}$ denotes the roll-out policy, and $v^{\mathrm{out}}_{t-1}$ is the corresponding soft value function. The roll-out policy is updated in a lazy manner—specifically, it is refreshed every $K$ steps using the current model parameters $p^{\theta_{s}}$. This lazy update scheme is critical for algorithmic stability in the offline setting, preventing rapid changes in the target while still allowing gradual improvement of the student policy toward higher reward. In the next section, we formalize this iterative distillation process as a complete algorithmic procedure.
@@ 4 Iterative Value-Weighted MLE
7:     Obtain a sample following the roll-out policy $p^{\mathrm{out}}_{t-1}$: $\bar{x}^{[i]}_{t-1}\sim p^{\mathrm{out}}_{t-1}(x^{[i]}_{t-1}\mid x^{[i]}_{t})$ for each $t$.
8:     Approximate soft-value functions $\hat{v}_{t-1}(\bar{x}^{[i]}_{t-1})\leftarrow r\left(\hat{x}_{0}(\bar{x}^{[i]}_{t-1};\theta^{\mathrm{out}})\right)$.
@@
$\theta_{s+1}\leftarrow\theta_{s}+\gamma\nabla_{\theta}\sum_{i=1}^{N}\sum_{t}\exp\left(\frac{\hat{v}_{t-1}(\bar{x}^{[i]}_{t-1})}{\alpha}\right)\log p_{t-1}^{\theta}(\bar{x}^{[i]}_{t-1}\mid x_{t}^{(i)}).$ (7)
@@ 4.1 Roll-In Phase
To balance these goals, we adopt a mixture roll-in strategy. At each training iteration, we sample the roll-in distribution from one of two candidate policies:
  • With probability $1-\beta_{s}$, we roll in from the pre-trained policy $p^{\text{pre}}_{t}$, which promotes exploration by generating diverse trajectories;
  • With probability $\beta_{s}$, we roll in from the current student policy $p^{\theta}_{t}$, which reflects the most up-to-date learned behavior and guides the model toward optimal behavior.
We gradually increase $\beta_{s}$ throughout training, allowing the roll-in distribution to transition from broad exploration in the early stages to reward-guided refinement in later stages. This adaptive schedule enables the model to progressively focus on high-reward regions as training progresses.
@@ 4.2 Roll-Out Phase
Recall that soft-value functions in (3) are defined as conditional expectation given $\bar{x}_{t-1}$. While one could estimate these expectations using Monte Carlo sampling or regression, as commonly done in standard RL settings, we recommend a more practical approximation: $\hat{v}_{t-1}(\bar{x}_{t-1}):=r\left(\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})\right)$. Recalling $\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})$ is the denoised prediction from the diffusion model parameterized by the current student policy. This approximation is based on the replacement of the expectation in (3) with its posterior mean.
@@ 6.1.1 Task Descriptors
We adopt EvoDiff (Alamdari et al., 2023) as our pre-trained diffusion model, a representative masked discrete diffusion model for protein sequence design, trained on the UniRef database (Suzek et al., 2007). We use EvoDiff as an unconditional generative model to produce natural protein sequences. To tackle downstream protein design tasks, we use the following reward functions inspired by prior work (Hie et al., 2022; Verkuil et al., 2022; Lisanza et al., 2024; Uehara et al., 2025b; Pacesa et al., 2024). Appendix D.1 provides detailed definitions of each reward function. For tasks involving secondary structure matching and globularity optimization, we employ ESMFold (Lin et al., 2023) to predict the 3D structures of generated sequences. For protein binder design, we leverage AlphaFold2 (Jumper et al., 2021) to model the 3D structure of the multimers. In order to encourage the naturalness and foldability of the designed proteins, we additionally incorporate structural confidence metrics, such as pLDDT and radius of gyration (Pacesa et al., 2024), as part of the reward function.
  • ss-match ($\beta$-sheet). This task aims to maximize the probability of secondary structure (SS) matching between the generated protein sequences and a predefined target pattern, computed across all residues. SS are predicted using the DSSP (Kabsch and Sander, 1983), and include $\alpha$-helices, $\beta$-sheets, and coils. Following Pacesa et al. (2024), we specifically encourage the formation of $\beta$-sheets, as protein generative models are known to exhibit a bias toward $\alpha$-helices.
  • Globularity. This task focuses on generating proteins with compact, globular 3D structures, a property known to enhance solubility and facilitate proper folding.
@@
Method Protein SS-match Protein Globularity DNA Enhancer HepG2 Molecule Docking - Parp1
$\beta$-sheet%↑ pLDDT↑ Globularity↓ pLDDT↑ Pred-Activity↑ ATAC-Acc↑ 3-mer Corr↑ Docking Score↑ NLL↓
Pre-trained 0.05 $\pm$ 0.05 0.37 $\pm$ 0.09 7.42 $\pm$ 2.59 0.37 $\pm$ 0.09 0.14 $\pm$ 0.26 $0$$\pm$ $0$0.15 7.2 $\pm$ 0.5 971 $\pm$ 32
Best-of-N (N=32) 0.26 $\pm$ 0.13 0.38 $\pm$ 0.11 5.05 $\pm$ 0.21 0.39 $\pm$ 0.14 1.30 $\pm$ 0.64 $0$$\pm$ $0$0.17 10.2 $\pm$ 0.4 951 $\pm$ 22
DRAKES - - - - 6.44 $\pm$ 0.04 0.825 $\pm$ 0.028 0.307 - -
Standard Fine-tuning 0.11 $\pm$ 0.12 0.32 $\pm$ 0.06 6.30 $\pm$ 1.11 0.31 $\pm$ 0.04 1.17 $\pm$ 1.23 0.094 $\pm$ 0.292 0.829 7.8 $\pm$ 1.8 908 $\pm$ 77
DDPP 0.60 $\pm$ 0.07 0.40 $\pm$ 0.08 5.98 $\pm$ 0.89 0.32 $\pm$ 0.10 5.18 $\pm$ 0.12 $0$$\pm$ 0.000 0.545 7.9 $\pm$ 1.3 981 $\pm$ 52
DDPO 0.19 $\pm$ 0.17 0.35 $\pm$ 0.10 4.75 $\pm$ 0.20 0.37 $\pm$ 0.06 7.09 $\pm$ 0.29 0.242 $\pm$ 0.428 0.283 8.5 $\pm$ 1.3 929 $\pm$ 43
VIDD 0.69 $\pm$ 0.13 0.69 $\pm$ 0.11 3.96 $\pm$ 0.16 0.56 $\pm$ 0.05 7.44 $\pm$ 0.34 0.898 $\pm$ 0.302 0.509 9.4 $\pm$ 1.7 741 $\pm$ 21
Table 2: Performance of different methods on protein binding design tasks w.r.t. ipTM, optimized reward, and diversity. The best result is highlighted in bold. To prevent model collapse, we report results with diversity $\geq$ 0.5.
@@
Figure 3: Protein structure visualizations for the PD-L1 and IFNAR2 binding design tasks. The target protein is shown in orange, with hotspot residues labeled on the structure.
(a) SS-Match
(b) Globularity
Figure 5: Sample efficiency and structural visualization results for the Secondary Structure Matching and Globularity tasks in protein generation. Panel (a) displays a $\beta$-barrel–like design, while panel (b) shows a TIM barrel–like design.
@@ 6.1.2 Baselines
  • Best-of-N: This is a naïve yet widely adopted approach for reward maximization at test time. Note that such inference-time methods are significantly slower compared to our fine-tuned models. Additional comparisons will be provided in Appendix E. We choose $N=32$ in general. Given that this method is 32 times slower than the baseline methods, we do not consider it a practical baseline for fine-tuning.
@@ 6.2 Results
(a) Molecules: Docking - Parp1
Figure 6: Generated samples from VIDD. For more samples, refer to Appendix F. Note that the surfaces and ribbons in (a) are representations of the target proteins, while the generated small molecules are displayed in the center.
@@
VIDD consistently outperforms baseline methods by achieving higher rewards in $\beta$-sheet content, globularity and ipTM. Figure 5 further demonstrates VIDD converges more rapidly under the same number of reward queries, indicating improved sample efficiency. Additionally, Figure 5 presents representative structures generated by VIDD, visually confirming the plausibility of the designed proteins. Figure 3 further illustrates the predicted binder-target complexes, confirming effective binding between the designed binders and target proteins.
@@
VIDD outperforms the fine-tuning baseline methods (DDPO, DDPP, and SFT) in terms of reward. Moreover, our method preserves a high degree of naturalness, as also visualized in Figure 6 by visualizing the generated samples. For more quantitative metrics that describe naturalness in molecules, refer to Appendix E.2.

### v2

@@ Abstract
We address the problem of fine-tuning diffusion models for reward-guided generation in biomolecular design. While diffusion models have proven highly effective in modeling complex, high-dimensional data distributions, real-world applications often demand more than high-fidelity generation, requiring optimization with respect to potentially non-differentiable reward functions such as physics-based simulation or rewards based on scientific knowledge. Although RL methods have been explored to fine-tune diffusion models for such objectives, they often suffer from instability, low sample efficiency, and mode collapse due to their on-policy nature. In this work, we propose an iterative distillation-based fine-tuning framework that enables diffusion models to optimize for arbitrary reward functions. Our method casts the problem as policy distillation: it collects off-policy data during the roll-in phase, simulates reward-based soft-optimal policies during roll-out, and updates the model by minimizing the KL divergence between the simulated soft-optimal policy and the current model policy. Our off-policy formulation, combined with KL divergence minimization, enhances training stability and sample efficiency compared to existing RL-based methods. Empirical results demonstrate the effectiveness and superior reward optimization of our approach across diverse tasks in protein, small molecule, and regulatory DNA design. The source code is released at (https://github.com/divelab/VIDD).
@@ 1 Introduction
Our contributions are summarized as follows. We propose a novel algorithm, VIDD, for fine-tuning diffusion models through iterative distillation of target policies composed of both value functions and the current policy. Unlike direct reward backpropagation methods, which require differentiable reward signals, our approach can handle non-differentiable rewards commonly encountered in scientific domains. Compared to PPO-based methods, our algorithm achieves more stable optimization. Algorithmically, our key innovation lies in effectively incorporating value functions specifically tailored to the diffusion models. Empirically, we validate the effectiveness of our method across a range of scientific tasks, particularly in protein and small molecular design.
@@ 1.1 Related Works
An alternative line of work focuses on training-free techniques that aim to improve rewards solely at inference time. The most straightforward example is Best-of-N sampling, which selects the highest-reward output from N generated candidates. More sophisticated methods go further by evaluating and selecting promising intermediate generation states during the sampling process (Wu et al., 2023; Li et al., 2024; Kim et al., 2025; Singhal et al., 2025; Ma et al., 2025). While these methods avoid the need for fine-tuning, they often incur significantly higher inference-time costs due to repeated sampling. Although they can boost reward in some cases, they do not directly improve the underlying generative model. In this respect, training-free methods are orthogonal and complementary to fine-tuning-based approaches, and the two can be effectively combined to achieve even stronger performance.
@@ 3.2 Iterative Distillation
$\mathop{\mathrm{argmin}}_{\theta}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}[\mathrm{KL}(p^{\star}_{t-1}(\cdot|x_{t})\|p^{\theta}_{t-1}(\cdot|x_{t}))],$ (4)
@@
$\mathop{\mathrm{argmax}}_{\theta}\sum_{t}\mathbb{E}_{x_{t}\sim u_{t}}\left[\frac{1}{\exp(v_{t}(x_{t})/\alpha)}\mathbb{E}_{x_{t-1}\sim p^{\mathrm{pre}}_{t-1}(x_{t-1}|x_{t})}[\exp(v_{t-1}(x_{t-1})/\alpha)\log p^{\theta}_{t-1}(x_{t-1}|x_{t}))]\right].$ (5)
@@
In practice, we adopt an iterative distillation procedure. This is motivated by the fact that, due to the reliance on empirical samples and approximate value functions, it is generally infeasible to accurately distill the target policy $p^{\star}_{t-1}$ into the student policy $p_{t-1}^{\theta}$ in a single step. Rather than fixing $p^{\mathrm{pre}}_{t-1}$ throughout training, we periodically update it by the latest student policy but in a lazy manner, allowing the target to gradually improve over time. This dynamic yet infrequent (lazy) target update enables progressive refinement, allowing both the teacher and student policies to evolve iteratively toward better alignment. This strategy is analogous to standard RL practices grounded in the Policy Improvement Theorem (Mnih et al., 2015).
@@
$\theta_{s+1}\leftarrow\theta_{s}+\nabla\sum_{t}\mathbb{E}_{x_{t}\sim u_{t},p^{\mathrm{out}}_{t-1}(x_{t-1}|x_{t})}\left[\frac{\exp(v^{\mathrm{out}}_{t-1}(x_{t-1})/\alpha)}{\exp(v_{t}(x_{t})/\alpha)}\log p^{\theta}_{t-1}(x_{t-1}|x_{t})\right].$ (6)
where $\{p^{\mathrm{out}}\}$ denotes the roll-out policy, and $v^{\mathrm{out}}_{t-1}$ is the corresponding soft value function. The roll-out policy is updated in a lazy manner—specifically, it is refreshed every $K$ steps using the current model parameters $p^{\theta_{s}}$. This lazy update scheme is critical for algorithmic stability in the offline setting, preventing rapid changes in the target while still allowing gradual improvement of the student policy toward higher reward. In the next section, we formalize this iterative distillation process as a complete algorithmic procedure.
@@ 4 Iterative Value-Weighted MLE
[…]
8:  Approximate soft-value functions $\hat{v}_{t-1}(\bar{x}^{[i]}_{t-1})\leftarrow r\left(\hat{x}_{0}(\bar{x}^{[i]}_{t-1};\theta^{\mathrm{out}})\right)$ and $\hat{v}_{t}(x^{[i]}_{t})\leftarrow r\left(\hat{x}_{0}(x^{[i]}_{t};\theta^{\mathrm{out}})\right)$.
@@
$\theta_{s+1}\leftarrow\theta_{s}+\gamma\nabla_{\theta}\sum_{i=1}^{N}\sum_{t}\left(\frac{\exp(\hat{v}_{t-1}(\bar{x}^{[i]}_{t-1})/\alpha)}{\exp(\hat{v}_{t}(x^{[i]}_{t})/\alpha)}\log p_{t-1}^{\theta}(\bar{x}^{[i]}_{t-1}\mid x_{t}^{(i)})\right).$ (7)
@@ 4.1 Roll-In Phase
To balance these goals, we adopt a mixture roll-in strategy by sampling the roll-in distribution from:
  • the pre-trained policy $p^{\text{pre}}_{t}$, which promotes exploration by generating diverse trajectories;
  • the roll-out policy $p^{\mathrm{out}}_{t}$ which is periodically updated and reflects stabilized knowledge from the student model.
At each training step, we sample from $p^{\text{pre}}_{t}$ with probability $1-\beta_{s}$, and from $p^{\mathrm{out}}_{t}$ with probability $\beta_{s}$. Further details on the construction of $p^{\mathrm{out}}_{t}$ are provided in the following subsection.
@@ 4.2 Roll-Out Phase
Recall that soft-value functions in (3) are defined as conditional expectation given $\bar{x}_{t-1}$. While one could estimate these expectations using Monte Carlo sampling or regression, as commonly done in standard RL settings, we recommend a more practical approximation: $\hat{v}_{t-1}(\bar{x}_{t-1}):=r\left(\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})\right)$. Recalling $\hat{x}_{0}(\bar{x}_{t-1};\theta^{\mathrm{out}})$ is the denoised prediction from the diffusion model parameterized by the current student policy. This approximation is based on the replacement of the expectation in (3) with its posterior mean. We apply the same approximation to $\hat{v}_{t}(x_{t})$, estimating it via $r(\hat{x}_{0}(x_{t};\theta^{\mathrm{out}}))$.
@@ 6.1.1 Task Descriptors
We adopt EvoDiff (Alamdari et al., 2023) as our pre-trained diffusion model, a representative masked discrete diffusion model for protein sequence design, trained on the UniRef database (Suzek et al., 2007). We use EvoDiff as an unconditional generative model to produce natural protein sequences. To tackle downstream protein design tasks, we use the following reward functions inspired by prior work (Hie et al., 2022; Verkuil et al., 2022; Lisanza et al., 2024; Uehara et al., 2025b; Pacesa et al., 2024). Appendix D.1 provides detailed definitions of each reward function. For tasks involving secondary structure matching optimization, we employ ESMFold (Lin et al., 2023) to predict the 3D structures of generated sequences. For protein binder design, we leverage AlphaFold2 (Jumper et al., 2021) to model the 3D structure of the multimers. In order to encourage the naturalness and foldability of the designed proteins, we additionally incorporate structural confidence metrics, such as pLDDT and radius of gyration (Pacesa et al., 2024), as part of the reward function.
  • $s$-match ($\beta$-sheet). This task aims to maximize the probability of secondary structure (SS) matching between the generated protein sequences and a predefined target pattern, computed across all residues. SS are predicted using the DSSP (Kabsch and Sander, 1983), and include $\alpha$-helices, $\beta$-sheets, and coils. Following Pacesa et al. (2024), we specifically encourage the formation of $\beta$-sheets, as protein generative models are known to exhibit a bias toward $\alpha$-helices.
@@
Method Protein SS-match DNA Enhancer HepG2 Molecule Docking - Parp1
$\beta$-sheet%↑ pLDDT↑ Diversity↑ Pred-Activity↑ ATAC-Acc↑ 3-mer Corr↑ Docking Score↑ NLL↓
Pre-trained 0.05 $\pm$ 0.05 0.37 $\pm$ 0.09 0.91 0.14 $\pm$ 0.26 0.000 $\pm$ 0.000 -0.15 7.2 $\pm$ 0.5 971 $\pm$ 32
Best-of-N (N=32) 0.26 $\pm$ 0.13 0.38 $\pm$ 0.11 0.90 1.30 $\pm$ 0.64 0.000 $\pm$ 0.000 -0.17 10.2 $\pm$ 0.4 951 $\pm$ 22
DRAKES - - - 6.44 $\pm$ 0.04 0.825 $\pm$ 0.028 0.307 - -
Standard Fine-tuning 0.48 $\pm$ 0.16 0.30 $\pm$ 0.04 0.57 1.17 $\pm$ 1.23 0.094 $\pm$ 0.292 0.829 7.8 $\pm$ 1.8 908 $\pm$ 77
DDPP 0.60 $\pm$ 0.07 0.40 $\pm$ 0.08 0.90 5.33 $\pm$ 0.94 0.305 $\pm$ 0.460 0.879 7.9 $\pm$ 1.3 981 $\pm$ 52
DDPO 0.81 $\pm$ 0.02 0.55 $\pm$ 0.05 0.52 7.38 $\pm$ 0.11 0.086 $\pm$ 0.280 0.398 8.5 $\pm$ 1.3 929 $\pm$ 43
VIDD 0.83 $\pm$ 0.01 0.82 $\pm$ 0.01 0.52 8.28 $\pm$ 0.18 0.820 $\pm$ 0.384 0.162 9.4 $\pm$ 1.7 741 $\pm$ 21
Table 2: Performance of different methods on protein binding design tasks w.r.t. ipTM, optimized reward, and diversity. The best result is highlighted in bold.
@@
Figure 3: Protein structure visualizations for the PD-L1 and IFNAR2 binding design tasks. The binder protein is shown in green and target protein is in orange, with hotspot residues labeled on the structure.
@@ 6.1.2 Baselines
  • Best-of-N: This is a naïve yet widely adopted approach for reward maximization at test time. Note that such inference-time methods are significantly slower compared to our fine-tuned models. Additional comparisons will be provided in Appendix E. Given that this method is $N$ times slower than the baseline methods, we do not consider it a practical baseline for fine-tuning.
@@ 6.1.3 Metrics
Figure 4: Protein structure visualizations for protein $S$-match tasks.
@@ 6.2 Results
VIDD consistently outperforms baseline methods by achieving higher rewards in $\beta$-sheet content and ipTM binding affinity. Additionally, Figure 3 and Figure 4 illustrate the predicted binder–target complexes and secondary-structure matching tasks, confirming effective binding of the designed binders to their targets and reasonable secondary structures of the designed proteins. Other quantitative results are provided in Appendix E.4, further validating the quality and effectiveness of the generated proteins.
@@
VIDD outperforms the fine-tuning baseline methods (DDPO, DDPP, and SFT) in terms of reward. For more quantitative metrics that describe naturalness in molecules, refer to Appendix E.3.